For each of the following numbers, find the smallest whole number by which it should be divided so as to get a perfect square. Also find the square root of the square number so obtained.   ii. 2925 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

2925 = 3×3×5×5×13
= (3×3)×(5×5)×13
Here, 13 cannot be paired.
∴ We will divide 2925 by 13 to get perfect square. New number = 2925 ÷ 13 = 225
square root of 225= 15
Thanks